a block of mass 20 kg hangs from a ceiling by a cord. calculate the tension in the cord


1
Expert's answer
2021-01-10T18:25:40-0500

Let's apply the Newton's Second Law of Motion:


"\sum F_y=ma_y=0,""T-mg=0,""T=mg=20\ kg\cdot 9.8\dfrac{m}{s^2}=196\ N."

Answer:

"T=196\ N."